# Provable Robustness in Multimodal Large Language Models via Feature Space Smoothing

### v1

@@ 5.4 Ablation study
An illustration of the implementation process is shown in Figure S.4.
S.9.2 Experimental results of using a lightweight U-net as noise purifier
The comparative results of different defense methods for image classification are summarized in Table S.6. In this experiment, the original diffusion-based noise purifier is replaced with a lightweight U-Net denoiser containing approximately 43.7M parameters. The results show that with this compact denoiser, FS-PSM consistently achieves strong and superior robustness, outperforming adversarial training–based baselines by a substantial margin. Furthermore, integrating FS-PSM with existing defenses yields additional robustness gains, leading to consistently favorable improvements across all evaluated settings. These results collectively demonstrate the effectiveness and generalizability of the proposed method.
S.9.3 Experimental results on large attack bound
Table S.7: Experimental results on adversarial robustness of different defense methods and MLLMs on image classification tasks. The attack bound $\epsilon=32/255$.
Model Method FOA [14]
@@
The experimental results under a large adversarial perturbation of $\varepsilon=32/255$ are presented in Table S.7. In addition, Figure S.5 provides visualizations of the adversarial examples generated by different attacks and perturbation bounds. These results show that nearly all existing defenses show great performance drop under such a strong attack. Under this challenging setting, the best accuracy is obtained by integrating FS-PSM with FARE, reaching 48.4% on LLaVA, while the lowest ASR (9.6%) is achieved by combining FS-PSM with TeCoA. These findings further highlight the effectiveness and adaptability of the proposed framework, even against large-magnitude adversarial perturbations.
Figure S.5: Visualization of the adversarial examples.
S.9.4 Analysis of robustness-efficiency trade-off
Table S.8: Analysis of the robustness-efficiency trade-off of using different $n_{0}$.
Model $n_{0}$ Robustness Efficiency
[…]
LLaVA- 1.5-7B (diffusion purifier) 1 0.595 86.2% 0.6% 0.28
2 0.652 88.4% 0.4% 0.41
4 0.697 88.8% 0.4% 0.66
8 0.719 89.6% 0.4% 1.12
16 0.728 89.2% 0.4% 2.10
LLaVA- 1.5-7B (lightweight purifier) 1 0.494 84.4% 0.2% 0.20
2 0.565 85.4% 0% 0.24
4 0.600 87.2% 0.2% 0.32
8 0.623 87.8% 0.4% 0.49
16 0.643 86.8% 0.4% 0.84
The robustness–efficiency trade-off results are summarized in Table S.8. In this experiment, we adopt the strong FOA attack as the adversary and report the average inference time per sample on LLaVA-1.5-7B using a single RTX 4090 GPU. The results show that increasing the sampling number $n_{0}$ consistently improves the FCS under adversarial perturbations. Specifically, increasing $n_{0}$ from 0 to 16 raises FCS from approximately 0.595 to 0.728, which aligns with our theoretical analysis in Section 3. However, this robustness gain comes at a substantial computational cost. For example, increasing $n_{0}$ from 4 to 16 raises the inference latency from 0.66 seconds to 2.10 seconds per sample, revealing an inherent trade-off between robustness and efficiency.

### v2

@@ A.4.1 Implementation details for main results
An illustration of the implementation process is shown in Figure A.4.
A.4.2 Experimental results on large attack bound
Table A.8: Experimental results on adversarial robustness of different defense methods and MLLMs on image classification tasks. The attack bound $\epsilon=32/255$.
Model Method FOA (Jia et al., 2025)
@@ A.4.2 Experimental results on large attack bound
The experimental results under a large adversarial perturbation of $\varepsilon=32/255$ are presented in Table A.8. In addition, Figure A.5 provides visualizations of the adversarial examples generated by different attacks and perturbation bounds. These results show that nearly all existing defenses show great performance drop under such a strong attack. Under this challenging setting, the best accuracy is obtained by integrating FS-GSB with FARE, reaching 48.4% on LLaVA, while the lowest ASR (9.6%) is achieved by combining FS-GSB with TeCoA. These findings further highlight the effectiveness and adaptability of the proposed framework, even against large-magnitude adversarial perturbations.
Figure A.5: Visualization of the adversarial examples.